# A New Construction Structure on Multi-access Coded Caching with Linear Subpacketization: Cyclic Multi-Access Non-Half-Sum Disjoint Packing

### v2

@@ V Performance analysis
In this section, we will present theoretical and numerical comparisons with the existing schemes [22, 25, 8, 23, 24, 28] respectively to show the performance of our new scheme in Theorem 4.
@@ V-A1 Comparison with the RK1 scheme in [22]
As a result, compared to the RK1 scheme in [22], our scheme either obtains smaller or same subpacketization. The multiplicative reduction amount by our transmission is $(2-\frac{2}{q})^{n}$. Only need $K\geq 3^{n}L$ ,the formula $(2-\frac{2}{q})^{n}\geq 1$ always holds.
@@ V-A2 Comparison with the CW scheme in [23]
As a result, compared to the CW scheme in [23], our scheme either obtains smaller subpacketization. The multiplicative reduction amount by our transmission is $\frac{2^{n}}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}$. Only when $K=qL$ ,the formula $\frac{2^{n}}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}=1$ always holds.
@@ V-A3 Comparison with the WCWL scheme in [25]
  • Otherwise, the WCWL scheme can be obtained as follow.
@@ V-A4 Comparison with the SR2 scheme in [28]
$\frac{F_{\text{SR2}}}{F}=\frac{q^{n}L}{q^{n}L}=1,\quad\frac{R_{\text{SR2}}}{R}=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}\left(2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}\right)}{2q^{n}}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}(L-1)+q^{n}}{2q^{n}L}>\frac{L-1}{2L}\left(2-\frac{6}{q}\right)^{n}.$
As a result, compared to the SR2 scheme in [28], our scheme either obtaining the same subpacketization. The multiplicative reduction amount by our transmission is at least $\frac{L-1}{2L}\left(2-\frac{6}{q}\right)^{n}$. When $q>6$ and $n\geq\frac{\ln\frac{2L}{L-1}}{\ln(2-\frac{6}{q})}$ ,the formula $\frac{L-1}{2L}\left(2-\frac{6}{q}\right)^{n}\geq 1$ always holds.
@@ V-B Numerical comparisons
Comparison with the schemes in [8, 22, 25, 23, 24]
When $K=84$ and $L=3$, we have our scheme and the schemes in [8, 22, 25, 23, 24] are shown in Fig. 3(a) and Fig. 3(b). It can be seen that our scheme can achieve a significant lower subpacketization and a slightly higher load than the schemes in [8, 23, 24, 22] with exponential subpaketization, a lower subpacketization and a same or lower load than the scheme in [25]. For instance, when $M/N=0.14$, we have our scheme with $F=84$ and $R=11.40$; the schemes in [8, 23, 22, 24] have lowest subpacketization is $F\approx 8.687\times{10}^{7}>84$ and lowest load is $R=4.10<11.40$. The WCWL scheme in [25] with $F\approx 170>84$ and $R\approx 27.98>11.40$.
(a) Memory ratio-subpacketization tradeoff for $K=84$, $L=3$
(b) Memory ratio-load tradeoff for $K=84$, $L=3$
Figure 3:
@@ VI Conclusion
This paper proposes a novel combinational structure called as cyclic multi-access non-half-sum disjoint packing (CMA-NHSDP), which is an extension of the traditional NHSDP structure to the MACC system. The CMA-NHSDP can realize a multi-access coded caching scheme with linear subpacketization while minimizing transmission load. In terms of performance comparison, compared to some existing schemes with linear subpacketization, our scheme has a larger coded caching gain, thus achieving a lower rate. And our scheme achieves far lower subpacketization and slightly higher load compared to some existing schemes with exponential subpacketization.
@@ Appendix A Proof of Theorem 3
Finally we check $L$-consecutive condition . All elements in $\mathscr{D}$ can be represented as $\sum_{j=i+1}^{n}\alpha_{j}\left(a_{j}x_{j}+g(j)\right)+\alpha_{i}\left(a_{i}x_{i}+g(i)\right)+\sum_{h=1}^{i-1}\alpha_{h}\left(a_{h}x_{h}+g(h)\right)$, where $\alpha_{i}\in\{-1,1\}$, $a_{i}\in[m_{i}]$. We consider the situation between neighboring elements in $\mathscr{D}$ after ascending sort:
@@
  • When $\alpha_{i}=-1$, $a_{i}=1$, and for any $h<i$, $\alpha_{h}=1$, $a_{h}=m_{h}$, i.e., $\sum_{j=i+1}^{n}\alpha_{j}\left(a_{j}x_{j}+g(j)\right)-\left(x_{i}+g(i)\right)+\sum_{h=1}^{i-1}\left(m_{h}x_{h}+g(h)\right)$, the neighboring larger element is $\sum_{j=i+1}^{n}\alpha_{j}\left(a_{j}x_{j}+g(j)\right)+\left(x_{i}+g(i)\right)-\sum_{h=1}^{i-1}\left(m_{h}x_{h}+g(h)\right)$, and the difference is $2\left(x_{i}+g(i)-\sum_{h=1}^{i-1}\left(f(h)+g(h)\right)\right)$; Let us consider the difference under modulo $L$
@@
    $\left\langle 2\left(x_{i}+g(i)-\sum_{j=1}^{i-1}\left(f(j)+g(j)\right)\right)\right\rangle_{L}=\left\langle 2\left(x_{1}+g(1)\right)\right\rangle_{L}=\left\langle 2\left(1+\frac{L-1}{2}\right)\right\rangle_{L}=1$ (22)
@@
    $\left\langle 2\left(x_{i}+g(i)-\sum_{j=1}^{i-1}\left(f(j)+g(j)\right)\right)\right\rangle_{L}$ $=\left\langle 2\left(2\sum_{j=1}^{i-1}\left(f(j)+g(j)\right)+1+g(i)-\sum_{j=1}^{i-1}\left(f(j)+g(j)\right)\right)\right\rangle_{L}$ (23)
    $=\left\langle 2\left(\sum_{j=1}^{i-1}\left(f(j)+g(j)\right)+1+g(i)\right)\right\rangle_{L}$
    $=\left\langle 2\left(\sum_{j=2}^{i}\left(f(j-1)+g(j)\right)+1+g(1)\right)\right\rangle_{L}$
    $=\left\langle 2\left(g(1)+1\right)\right\rangle_{L}$
    $=1$

### v3

@@ V Performance analysis
In this section, we will present theoretical and numerical comparisons with the existing schemes [22, 25, 8, 23, 24, 28, 26, 27, 29, 31] respectively to show the performance of our new scheme in Theorem 4.
@@ V-A1 Comparison with the RK1 scheme in [22]
As a result, compared to the RK1 scheme in [22], our scheme either obtaining smaller or same subpacketization. The multiplicative reduction amount by our transmission is $(2-\frac{2}{q})^{n}$. Only when $K\geq 3^{n}L$ ,the formula $(2-\frac{2}{q})^{n}\geq 1$ always holds.
@@ V-A2 Comparison with the CW scheme in [23]
As a result, compared to the CW scheme in [23], our scheme either obtaining smaller subpacketization. The multiplicative reduction amount by our transmission is $\frac{2^{n}}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}$. Only when $K=qL$ ,the formula $\frac{2^{n}}{q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}+1}=1$ always holds.
@@ V-A3 Comparison with the WCWL scheme in [25]
  • Otherwise, the WCWL scheme can be obtain as follow.
@@ V-A4 Comparison with the SR2 scheme in [28]
$\frac{F_{\text{SR2}}}{F}=\frac{q^{n}L}{q^{n}L}=1,\quad\frac{R_{\text{SR2}}}{R}=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}\left(2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}\right)}{2q^{n}}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}(L-1)+q^{n}}{2q^{n}L}>\frac{1}{2L}\left(2-\frac{6}{q}\right).$
As a result, compared to the SR2 scheme in [28], our scheme either obtaining the same subpacketization. The multiplicative reduction amount by our transmission is at least $\frac{1}{2L}\left(2-\frac{6}{q}\right)$. When $q>6$ and $n\geq\frac{\ln 2L}{\ln(2-\frac{6}{q})}$ ,the formula $\frac{1}{2L}\left(2-\frac{6}{q}\right)\geq 1$ always holds.
@@ V-B Numerical comparisons
V-B1 Comparison with the schemes in [8, 22, 25, 23, 24]
When $K=84$ and $L=3$, we have our scheme and the schemes in [8, 22, 25, 23, 24] are shown in Fig. 4 and Fig. 4. It can be seen that our scheme can achieve a significant lower subpacketization and a slightly higher load than the schemes in [8, 23, 24, 22] with exponential subpaketization, a lower subpacketization and a same or lower load than the scheme in [25]. For instance, when $M/N=0.14$, we have our scheme with $F=84$ and $R=11.40$; the schemes in [8, 23, 22, 24] have lowest subpacketization is $F\approx 8.687\times{10}^{7}>84$ and lowest load is $R=4.10<11.40$. The WCWL scheme in [25] with $F\approx 170>84$ and $R\approx 27.98>11.40$.
Figure 3: Memory ratio-subpacketization tradeoff for $K=84$, $L=3$
Figure 4: Memory ratio-load tradeoff for $K=84$, $L=3$
V-B2 Comparison with the schemes in  [26, 27, 28, 29, 31]
We compared our proposed scheme in Theorem 4 with the schemes in [26], [27], [28], [29], and [31] by selecting specific user numbers and memory ratios, as listed in Table II. Compared with the SPE scheme in [26], under similar numbers of users and memory ratios, our scheme achieves significantly lower subpacketization at the cost of a slightly higher transmission load. Compared with the SR1 scheme in [27], for the same number of users and memory ratio, the proposed scheme achieves a significantly lower transmission load while maintaining smaller subpacketization. For example, when $M/N=0.0222$ and $K=45$, the transmission load is reduced from $20$ to $5$, and the subpacketization is reduced from $90$ to $45$. Compared with the SR2 scheme in [28], the proposed scheme achieves a lower transmission load and the same subpacketization level while having a close number of users and a close memory ratio. For example, when $K=576$ and $M/N\approx 0.063$, the load decreases from $63$ to $9$ while keeping the same subpacketization. Compared with the MR scheme in [29] and the NR scheme in [31], for the same number of users and memory ratio, the proposed scheme always achieves a lower or equal transmission load while maintaining the same subpacketization. For example, when $M/N=0.1852$ and $K=27$, the proposed scheme achieves a load of $R=3$, which is lower than the $R=5.3333$ achieved by the NR scheme.
TABLE II: The numerical comparison between the scheme in Theorem 4 and the schemes in [26, 27, 28, 29, 31]
$K$ $M/N$ Scheme Parameters Load Subpacketization
$28$ $0.0714$ SPE scheme in [26] $z=3$ $6.7500$ $168$
$27$ $0.0370$ Scheme in Theorem 4 $(v,n,L)=(27,1,3)$ $12$ $27$
$64$ $0.0313$ SPE scheme in [26] $z=7$ $14.8846$ $832$
$63$ $0.0159$ Scheme in Theorem 4 $(v,n,L)=(63,1,7)$ $28$ $63$
$288$ $0.0069$ SPE scheme in [26] $z=7$ $89.3841$ $19872$
$287$ $0.0035$ Scheme in Theorem 4 $(v,n,L)=(287,1,7)$ $140$ $287$
$1112$ $0.0018$ SPE scheme in [26] $z=11$ $360.0311$ $303576$
$1111$ $0.0009$ Scheme in Theorem 4 $(v,n,L)=(1111,1,11)$ $550$ $1111$
$45$ $0.0222$ SR1 scheme in [27] $(k,z)=(1,5)$ 2020 $90$
$45$ $0.0222$ Scheme in Theorem 4 $(v,n,L)=(45,2,5)$ 55 $45$
$147$ $0.0884$ SR1 scheme in [27] $(k,z)=(13,3)$ $54$ $294$
$147$ $0.0884$ Scheme in Theorem 4 $(v,n,L)=(147,2,3)$ $27$ $147$
$245$ $0.0531$ SR1 scheme in [27] $(k,z)=(13,5)$ $90$ $490$
$245$ $0.0531$ Scheme in Theorem 4 $(v,n,L)=(245,2,5)$ $45$ $245$
$1331$ $0.0158$ SR1 scheme in [27] $(k,z)=(21,11)$ $550$ $2662$
$1331$ $0.0158$ Scheme in Theorem 4 $(v,n,L)=(1331,2,11)$ $275$ $1331$
$80$ $0.1250$ SR2 scheme in [28] $(k,L)=(10,5)$ $7.5$ $80$
$80$ $0.0875$ Scheme in Theorem 4 $(v,n,L)=(80,2,5)$ 55 $80$
$144$ $0.0625$ SR2 scheme in [28] $(k,L)=(9,9)$ $15.75$ $144$
$144$ $0.0486$ Scheme in Theorem 4 $(v,n,L)=(144,2,9)$ 99 $144$
$576$ $0.0625$ SR2 scheme in [28] $(k,L)=(36,9)$ $63$ $576$
$576$ $0.0642$ Scheme in Theorem 4 $(v,n,L)=(576,3,9)$ 99 $576$
$996$ $0.1667$ SR2 scheme in [28] $(k,L)=(166,5)$ $27.6667$ $996$
$1215$ $0.1737$ Scheme in Theorem 4 $(v,n,L)=(1215,5,5)$ 55 $1215$
$27$ $0.1852$ MR scheme in [29] $(i,L)=(5,3)$ 33 $27$
$27$ $0.1852$ Scheme in Theorem 4 $(v,n,L)=(27,2,3)$ 33 $27$
$45$ $0.0222$ MR scheme in [29] $(i,L)=(1,5)$ 2020 $45$
$45$ $0.0222$ Scheme in Theorem 4 $(v,n,L)=(45,2,5)$ 55 $45$
$245$ $0.0531$ MR scheme in [29] $(i,L)=(13,5)$ $90$ $245$
$245$ $0.0531$ Scheme in Theorem 4 $(v,n,L)=(245,2,5)$ $45$ $245$
$3125$ $0.1181$ MR scheme in [29] $(i,L)=(369,5)$ $320$ $3125$
$3125$ $0.1181$ Scheme in Theorem 4 $(v,n,L)=(3125,4,5)$ $80$ $3125$
$27$ $0.1852$ NR scheme in [31] $(i,L)=(5,3)$ $5.3333$ $27$
$27$ $0.1852$ Scheme in Theorem 4 $(v,n,L)=(27,2,3)$ 33 $27$
$45$ $0.0222$ NR scheme in [31] $(i,L)=(1,5)$ $35.5556$ $45$
$45$ $0.0222$ Scheme in Theorem 4 $(v,n,L)=(45,2,5)$ 55 $45$
$245$ $0.0531$ NR scheme in [31] $(i,L)=(13,5)$ $132.2449$ $245$
$245$ $0.0531$ Scheme in Theorem 4 $(v,n,L)=(245,2,5)$ $45$ $245$
$1331$ $0.0158$ NR scheme in [31] $(i,L)=(21,11)$ $909.0909$ $1331$
$1331$ $0.0158$ Scheme in Theorem 4 $(v,n,L)=(1331,2,11)$ $275$ $1331$
@@ VI Conclusion
This paper proposes a novel combinational structure called as cyclic multi-access non-half-sum disjoint packing (CMA-NHSDP), which is an extension of the traditional NHSDP structure to MACC system. The CMA-NHSDP can realize a multi-access coded caching scheme with linear subpacketization while minimizing transmission load. In terms of performance comparison, compared to some existing schemes with linear subpacketization, our scheme has larger coded caching gain, thus achieving a lower transmission load. And our scheme achieves far lower subpacketization and slightly higher load compare to some existing schemes with exponential subpacketization.
@@ Appendix A Proof of Theorem 3
Finally, we verify the $L$-consecutive condition. Recall that every element in $\mathscr{D}$ can be represented as $\sum_{j=i+1}^{n}\alpha_{j}\left(a_{j}x_{j}+g(j)\right)+\alpha_{i}\left(a_{i}x_{i}+g(i)\right)+\sum_{h=1}^{i-1}\alpha_{h}\left(a_{h}x_{h}+g(h)\right)$, where $i\in[n]$, $\alpha_{i}\in\{-1,1\}$, $a_{i}\in[m_{i}]$. To show that $\mathbb{Z}_{v}\setminus\mathcal{D}$ can be partitioned into disjoint subsets of $L$ consecutive integers, it suffices to prove that for any two neighboring elements in the sorted order of $\mathcal{D}$, their difference is equal to $1$ modulo $L$. We analyze the possible patterns of the coefficients. Due to the recursive construction of $\mathcal{D}$, the neighboring elements can be categorized into three cases.
@@
  • Case 2: When $\alpha_{i}=-1$, $a_{i}=1$, $\alpha_{h}=1$, and $a_{h}=m_{h}$, i.e., $\sum_{j=i+1}^{n}\alpha_{j}\left(a_{j}x_{j}+g(j)\right)-\left(x_{i}+g(i)\right)+\sum_{h=1}^{i-1}\left(m_{h}x_{h}+g(h)\right)$, the neighboring larger element is $\sum_{j=i+1}^{n}\alpha_{j}\left(a_{j}x_{j}+g(j)\right)+\left(x_{i}+g(i)\right)-\sum_{h=1}^{i-1}\left(m_{h}x_{h}+g(h)\right)$, and the difference is $2\left(x_{i}+g(i)-\sum_{h=1}^{i-1}\left(m_{h}x_{h}+g(h)\right)\right)$; Let us consider the difference under modulo $L$
@@
    $\left\langle 2\left(x_{i}+g(i)-\sum_{h=1}^{i-1}\left(m_{h}x_{h}+g(h)\right)\right)\right\rangle_{L}=\left\langle 2\left(x_{1}+g(1)\right)\right\rangle_{L}=\left\langle 2\left(1+\frac{L-1}{2}\right)\right\rangle_{L}=1.$ (22)
@@
    $\left\langle 2\left(x_{i}+g(i)-\sum_{h=1}^{i-1}\left(m_{h}x_{h}+g(h)\right)\right)\right\rangle_{L}$ $=\left\langle 2\left(2\sum_{h=1}^{i-1}\left(m_{h}x_{h}+g(h)\right)+1+g(i)-\sum_{h=1}^{i-1}\left(m_{h}x_{h}+g(h)\right)\right)\right\rangle_{L}$ (23)
    $=\left\langle 2\left(\sum_{h=1}^{i-1}\left(m_{h}x_{h}+g(h)\right)+1+g(i)\right)\right\rangle_{L}$
    $=\left\langle 2\left(\sum_{h=2}^{i}\left(m_{h-1}x_{h-1}+g(h)\right)+1+g(1)\right)\right\rangle_{L}$
    $=\left\langle 2\left(g(1)+1\right)\right\rangle_{L}=1.$
[…]